OpenStudy (anonymous):

There are 15 rows of diamond rings in a case in the jewelry store. There are 24 rings in each row. Another case has 12 rows of rings with a variety of colored stones and has 18 rings in each row. How many more diamond rings than rings with colored stones are in the jewelry store?

OpenStudy (anonymous):

Please answer

OpenStudy (texaschic101):

15 rows of diamond rings...24 rings in each row 15 * 24 = 360 diamond rings 12 rows of colored stones...18 rings in each row 12 * 18 = 216 colored rings Leaving a total of (360 - 216) = 144 more diamonds then colored stones